OpenStudy (anonymous):

Solve the system by graphing. -2x= 2y - 4 2x - y = -5

OpenStudy (anonymous):

I am so not good at graphing. I just don't get this.

OpenStudy (australopithecus):

x is your input y is your output First isolate the y for both of these equations

OpenStudy (australopithecus):

once you have done that I will tell you the next step

OpenStudy (anonymous):

ok

OpenStudy (australopithecus):

by isolate I mean move y to one side of the equation and the rest of the junk to the other side of the equation

OpenStudy (australopithecus):

If you get stuck with the algebra I can help you just show me where you get stuck

OpenStudy (anonymous):

ok let me see if Im doing this right. when I isolate y for the first equation of -2x = 2y - 4, would it be 2y = -2x + 4?

OpenStudy (australopithecus):

yes that is the first step 2y = -2x + 4 now to isolate the y further we need to eliminate the 2, we can do this by dividing both sides of the equation by 2, remember that 2/2 = 1 and y*1 = y so 2y/2 = -2x/2 + 4/2 1(y) = (-2x + 4)/2 y = (-2x + 4)/2

OpenStudy (australopithecus):

Now just do the same for the next equation and I will tell you what to do next

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x - y = -5 convert to: y = 2x + 5?

OpenStudy (australopithecus):

looks good :)

OpenStudy (anonymous):

ok so now I have y = -2x + 2 and y = 2x + 5 the x's cancel out making y = 7?

OpenStudy (australopithecus):

ok so we have: y = 2x + 5 y = -2x/2 + 4/2 both equations are linear so first find the x and y intercepts you find the x-intercept by setting y = 0 so you get point (x,0) to find the y-intercept you set x = 0 so you get a point (0,y) next just graph a variety of points, by inserting values into x and seeing what you get out as y so for example graph -10, -5, -4, -2, 0, 2, 4, 5, 10 what this problem is asking you is where these two equations intersect

OpenStudy (anonymous):

okay now Im lost lol Do I graph just points on x. Graphing is where I am completely confused

OpenStudy (australopithecus):

 using this axis, each time you insert a value into x you get a value y, this gives you a point on the graph (x,y) plot the points and connect them with a line

OpenStudy (australopithecus):

so for the second equation y = -x + 2 I'm going to find the intercepts to find the x-intercept I set y=0, giving me a point (x,0) 0 = -x + 2 -2 = -x 2 = x so I have a x-intercept of (2,0) now I'm going to find the y-intercept by setting x = 0 giving me a point (0,y) y = -0 + 2 y = 2 so I have a y-intercept of (0,2) I'm going to insert -10 into the x value to give me point (-10,y) y = -(-10) + 2 y = 10 + 2 y = 12 so I have point (-10,12) I'm going to insert -5 into the x value to give me point (-5,y) y = -(-5) + 2 y = 5 + 2 y = 7 so I have a point (-5, 7) Now I graph the points I have and connect them with a line

OpenStudy (anonymous):

ok

OpenStudy (australopithecus):